OpenStudy (anonymous):

The sum of the interior angles, s, in an n-sided polygon can be determined using the formula s = 180(n – 2), where n is the number of sides. Benita solves this equation for n and writes the equivalent equation n = s/180 + 2. Using this formula, how many sides does a polygon have if the sum of the interior angles is 1,260°? _______ sides

OpenStudy (anonymous):

plug 1260 for s in the equation and solve for N use n= s/180+2 = 1260/180 + 2= 7 +2 =9

OpenStudy (anonymous):

n=9